# Weak-to-Strong Diffusion with Reflection

### v2

@@ Abstract
The goal of diffusion generative models is to align the learned distribution with the real data distribution through gradient score matching. However, inherent limitations in training data quality, modeling strategies, and architectural design lead to inevitable gap between generated outputs and real data. To reduce this gap, we propose Weak-to-Strong Diffusion (W2SD), a novel framework that utilizes the estimated difference between existing weak and strong models (i.e., weak-to-strong difference) to approximate the gap between an ideal model and a strong model. By employing a reflective operation that alternates between denoising and inversion with weak-to-strong difference, we theoretically understand that W2SD steers latent variables along sampling trajectories toward regions of the real data distribution. W2SD is highly flexible and broadly applicable, enabling diverse improvements through the strategic selection of weak-to-strong model pairs (e.g., DreamShaper vs. SD1.5, good experts vs. bad experts in MoE). Extensive experiments demonstrate that W2SD significantly improves human preference, aesthetic quality, and prompt adherence, achieving SOTA performance across various modalities (e.g., image, video), architectures (e.g., UNet-based, DiT-based, MoE), and benchmarks. For example, Juggernaut-XL with W2SD can improve with the HPSv2 winning rate up to 90% over the original results. Moreover, the performance gains achieved by W2SD markedly outweigh its additional computational overhead, while the cumulative improvements from different weak-to-strong difference further solidify its practical utility and deployability. The code is publicly available at github.com/xie-lab-ml//Weak-to-Strong-Diffusion-with-Reflection.
@@ 1 Introduction
In probabilistic modeling, the estimated density represents model-predicted distributions while the ground truth density reflects actual data distributions. Diffusion models (Song et al., ), known for its powerful generative capabilities and diversity, estimate the gradient of the probability density in perturbed data distributions by optimizing a score-based network, which has become a mainstream paradigm in many generative tasks (Podell et al., ; Guo et al., 2023).
Figure 1: W2SD leverages the gap between weak and strong models to approximate the gap between strong and ideal models.
@@
In this work, we demonstrate that the gap between the strong and weak models can be used to empirically approximate the gap between the ideal and strong models. In Figure 1, we propose to use the estimated weak-to-strong difference $\Delta_{1}$ to approximate the inaccessible strong-to-ideal difference $\Delta_{2}$, thereby enhancing the strong model toward the ideal model. We note this weak-to-strong concept has been widely applied in training processes of machine learning, with early instances such as AdaBoost (Schapire, 2013), where the ensembling of weak models enhances the performance of strong models. Also, Burns et al. (2023) have demonstrated that weak models can serve as supervisory signals to assist in the alignment during the training of large LLMs.
To empirically estimate the weak-to-strong difference, we draw upon reflective mechanisms, the process of modifying generated outputs based on prior states, which have been extensively studied in the field of LLMs (Gou et al., ; Madaan et al., 2024; Shinn et al., 2024). Specifically, we propose Weak-to-Strong Diffusion (W2SD), a novel framework that leverages weak-to-strong difference to approximate the strong-to-ideal difference. By incorporating a reflective operation that alternates between denoising and inversion based on the weak-to-strong difference, we theoretically understand in Section 3.1 that W2SD guides latent variables along sampling trajectories, effectively steering them toward regions of the real data distribution. And we provide qualitative results in Figure 2
@@
First, we introduce the weak-to-strong concept into the inference enhancement of diffusion models, demonstrating that the difference of estimated densities between weak and strong diffusion models can approximate the difference between strong and ideal models, consequently bridging the gap between the learned and real data distribution.
@@ 3.1 Weak-to-Strong Diffusion
In diffusion models, the goal is to minimize the difference between the estimated density gradient and ground truth density gradients. However, due to the real data distribution is inaccessible, this difference cannot be directly quantified. To tackle this issue, we consider models of varying capacities: a strong model (with its corresponding denoising process denoted as $\mathcal{M}^{\mathrm{s}}$ and the estimated density as $p^{\mathrm{s}}$) and a weak model (similarly, $\mathcal{M}^{\mathrm{w}}$ and $p^{\mathrm{w}}$).
As shown in  Figure 1, we define the weak-to-strong difference as $\Delta_{1}=\nabla p^{\mathrm{s}}-\nabla p^{\mathrm{w}}$, and the strong-to-ideal difference as $\Delta_{2}=\nabla p^{\mathrm{gt}}-\nabla p^{\mathrm{s}}$. By approximating $\Delta_{2}$ using the estimable $\Delta_{1}$, we indirectly reduce the gap between existing diffusion models and the ideal model, bringing the learned distribution closer to the real data distribution. In the next subsection, we introduce how to achieve this approximation from a reflective perspective.
@@
Suppose $x_{t}$ is the latent variable at time $t$, let $p_{t}^{\mathrm{s}}$ and $p_{t}^{\mathrm{w}}$ denote the probability density estimates derived from $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$, respectively. The reflective operator $\mathcal{M}^{\mathrm{w}}_{\mathrm{inv}}(\mathcal{M}^{\mathrm{s}}(\cdot))$ refines $x_{t}$ to $\tilde{x}_{t}$ as
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}p_{t}^{\mathrm{s}}(x_{t})-\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})),$ (6)
where $\Delta_{1}(t)=\nabla_{x_{t}}p_{t}^{\mathrm{s}}(x_{t})-\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})$ represents the weak-to-strong difference between $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$ at time $t$.
@@ 2-D Gaussian Mixture Data
We also visualize the mechanism of W2SD on 2D scenario. In Figure 5 (first row), we modulate the proportion of the training data to obtain $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$. W2SD balances the distribution by exploiting the discrepancy between $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$ (the region in the bottom-left corner, denoted as $\Delta_{1}$) to approximate the unattainable strong-to-ideal difference $\Delta_{2}$, and increase the sampling probability for the bottom-left region. In Figure 5 (second row), We visualize the denoising trajectories under different settings, further validating the effectiveness of W2SD.
@@ 6 Conclusion
To the best of our knowledge, this work is the first to systematically integrate the weak-to-strong mechanism into the inference enhancement of diffusion models. We theoretically and empirically understand that the estimated weak-to-strong difference can effectively approximate the strong-to-ideal difference, enhancing the alignment between the learned distributions from existing diffusion models and the real data distribution. Building on this concept, we propose W2SD, which utilizes the estimated difference in density gradients to optimize sampling trajectories via reflective operations. W2SD demonstrates its effectiveness as a general-purpose framework through its cumulative performance improvements, flexible definition of weak-to-strong model pairs, and efficient performance gains with minimal computational overhead. We hope this work inspires interest in diffusion model scaling laws, advances sampling methods, and deepens understanding of probabilistic modeling in generative models.
@@ Appendix A Related Work
The concept of improving weak models into strong models originates from the AdaBoost (Høgsgaard et al., 2023), which constructs a more accurate classifier by aggregating multiple weak classifiers. Building upon this,  Green Larsen & Ritzert (2022); Høgsgaard et al. (2023) introduced a provably optimal weak-to-strong learner, establishing a robust theoretical foundation for this weak-to strong paradigm. In the field of LLMs, several studies (Chen et al., ; Burns et al., 2023) have utilized weak models as supervisory signals to facilitate the alignment of LLMs. This paradigm of weak-to-strong generation during training has similarly been investigated in the context of diffusion model training (Chen et al., 2025). And Sugiyama et al. (2013) recommended directly estimating the density difference between weak and strong models instead of separate estimations. In this work, we propose W2SD, a general framework that, for the first time, implements weak-to-strong enhancement in diffusion inference through a reflective mechanism.
@@ D.1 Proof of Theorem 1
where $s_{\theta}^{\mathrm{s}}$ denotes the score predicted by $\mathcal{M}^{\mathrm{s}}$, which is equivalent to the gradient of the probability distribution $p_{t}^{\mathrm{s}}$, so, Equation 7 can be rewritten as
$x_{t-\Delta t}=x_{t}+\sigma^{2t}\nabla_{x_{t}}p_{t}^{\mathrm{s}}(x_{t})\Delta t.$ (8)
@@
where $s_{\theta}^{\mathrm{w}}$ denotes the score predicted by $\mathcal{M}^{\mathrm{w}}$. Since $\Delta t$ is typically small, we neglect the approximation error in the diffusion inversion process which implies that $s_{\theta}^{\mathrm{w}}(x_{t-\Delta t},t)$ is equivalent to the gradient of the probability distribution $p_{t}^{\mathrm{w}}$. Hence we can reformulate Equation 9 as
$\tilde{x}_{t}=x_{t-\Delta t}-\sigma^{2t}\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})\Delta t.$ (10)
@@
$\tilde{x}_{t}$ $=x_{t-\Delta t}-\sigma^{2t}\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})\Delta t$ (11)
$=x_{t}+\sigma^{2t}\nabla_{x_{t}}p_{t}^{s}(x_{t})\Delta t-\sigma^{2t}\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})\Delta t$ (12)
$=x_{t}+\sigma^{2t}\Delta t(\underbrace{\nabla_{x_{t}}p_{t}^{s}(x_{t})-\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})}_{\text{weak-to-strong difference}\Delta_{1}(t)}).$ (13)
From Equation 13, we observe that for the latent variable $x_{t}$, the reflection operator $\mathcal{M}^{\mathrm{w}}_{\mathrm{inv}}(\mathcal{M}^{\mathrm{s}}(\cdot))$ in W2SD perturbs $x_{t}$ along the direction of $\Delta_{1}(t)$, producing the refined variable $\tilde{x}_{t}$. When the weak-to-strong difference $\Delta_{1}$ closely approximates the unattainable strong-to-ideal difference $\Delta_{2}(t)$, the resulting $\tilde{x}_{t}$ becomes more aligned with the ground truth ideal distribution.
@@ E.1 The Impact of Approximation Error in the Inversion Process
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}p_{t}^{\mathrm{s}}(x_{t})-\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})-\textrm{E}_{t}),$ (16)
@@
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}p_{t}^{\mathrm{s}}(x_{t})-\nabla_{x_{t}}p_{t}^{\mathrm{w}}(x_{t})-k\epsilon).$ (17)
@@
E.2 Relationship with Re-Sampling

### v3

@@ Abstract
The goal of diffusion generative models is to align the learned distribution with the real data distribution through gradient score matching. However, inherent limitations in training data quality, modeling strategies, and architectural design lead to inevitable gap between generated outputs and real data. To reduce this gap, we propose Weak-to-Strong Diffusion (W2SD), a novel framework that utilizes the estimated difference between existing weak and strong models (i.e., weak-to-strong difference) to bridge the gap between an ideal model and a strong model. By employing a reflective operation that alternates between denoising and inversion with weak-to-strong difference, we theoretically understand that W2SD steers latent variables along sampling trajectories toward regions of the real data distribution. W2SD is highly flexible and broadly applicable, enabling diverse improvements through the strategic selection of weak-to-strong model pairs (e.g., DreamShaper vs. SD1.5, good experts vs. bad experts in MoE). Extensive experiments demonstrate that W2SD significantly improves human preference, aesthetic quality, and prompt adherence, achieving SOTA performance across various modalities (e.g., image, video), architectures (e.g., UNet-based, DiT-based, MoE), and benchmarks. For example, Juggernaut-XL with W2SD can improve with the HPSv2 winning rate up to 90% over the original results. Moreover, the performance gains achieved by W2SD markedly outweigh its additional computational overhead, while the cumulative improvements from different weak-to-strong difference further solidify its practical utility and deployability. The code is publicly available at github.com/xie-lab-ml/Weak-to-Strong-Diffusion-with-Reflection.
@@ 1 Introduction
In probabilistic modeling, the estimated density represents model-predicted distributions while the ground truth density reflects actual data distributions. Diffusion models (Song et al., ), known for its powerful generative capabilities and diversity, estimate the gradient of the log probability densities in perturbed data distributions by optimizing a score-based network, which has become a mainstream paradigm in many generative tasks (Podell et al., ; Guo et al., 2023).
Figure 1: W2SD leverages the gap between weak and strong models to bridge the gap between strong and ideal models.
@@
In this work, we demonstrate that the gap between the strong and weak models can be used to empirically bridge the gap between the ideal and strong models. In Figure 1, we propose to use the estimated weak-to-strong difference $\Delta_{1}$ to bridge the inaccessible strong-to-ideal difference $\Delta_{2}$, thereby enhancing the strong model toward the ideal model. Further evidence supporting the viability of using $\Delta_{1}$ to bridge $\Delta_{2}$ is provided in Section E.1.
We note this weak-to-strong concept has been widely applied in training processes of machine learning, with early instances such as AdaBoost (Schapire, 2013), where the ensembling of weak models enhances the performance of strong models. Also, Burns et al. (2023) have demonstrated that weak models can serve as supervisory signals to assist in the alignment during the training of large LLMs.
To empirically estimate the weak-to-strong difference, we draw upon reflective mechanisms, the process of modifying generated outputs based on prior states, which have been extensively studied in the field of LLMs (Gou et al., ; Madaan et al., 2024; Shinn et al., 2024). Specifically, we propose Weak-to-Strong Diffusion (W2SD), a novel framework that leverages weak-to-strong difference to bridge the strong-to-ideal difference. By incorporating a reflective operation that alternates between denoising and inversion based on the weak-to-strong difference, we theoretically understand in Section 3.1 that W2SD guides latent variables along sampling trajectories, effectively steering them toward regions of the real data distribution. And we provide qualitative results in Figure 2.
@@
First, we introduce the weak-to-strong concept into the inference enhancement of diffusion models, demonstrating that the gradient difference of estimated log probability densities between weak and strong diffusion models can approximate the difference between strong and ideal models, consequently bridging the gap between the learned and real data distribution.
@@ 3.1 Weak-to-Strong Diffusion
In diffusion models, the goal is to minimize the gradients difference of log probability densities between estimated results and ground truth. However, due to the real data distribution is inaccessible, this difference cannot be directly quantified. To tackle this issue, we consider models of varying capacities: a strong model (with its corresponding denoising process denoted as $\mathcal{M}^{\mathrm{s}}$ and the estimated density as $p^{\mathrm{s}}$) and a weak model (similarly, $\mathcal{M}^{\mathrm{w}}$ and $p^{\mathrm{w}}$).
As shown in  Figure 1, we define the weak-to-strong difference as $\Delta_{1}=\nabla\log{p^{\mathrm{s}}}-\nabla\log{p^{\mathrm{w}}}$ and the strong-to-ideal difference as $\Delta_{2}=\nabla\log{p^{\mathrm{gt}}}-\nabla\log{p^{\mathrm{s}}}$. By approximating $\Delta_{2}$ using the estimable $\Delta_{1}$, we indirectly reduce the gap between existing diffusion models and the ideal model, bringing the learned distribution closer to the real data distribution. In the next subsection, we introduce how to achieve this approximation from a reflective perspective.
@@
Suppose $x_{t}$ is the latent variable at time $t$, let $p_{t}^{\mathrm{s}}$ and $p_{t}^{\mathrm{w}}$ denote the probability density estimates derived from $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$. The reflective operator $\mathcal{M}^{\mathrm{w}}_{\mathrm{inv}}(\mathcal{M}^{\mathrm{s}}(\cdot))$ refines $x_{t}$ to $\tilde{x}_{t}$ as
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}\log{p_{t}^{\mathrm{s}}(x_{t})}-\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}),$ (6)
where $\Delta_{1}(t)=\nabla_{x_{t}}\log{p_{t}^{\mathrm{s}}(x_{t})}-\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}$ means the weak-to-strong difference between $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$ at time $t$.
It is noted that Karras et al. (2024) proposed Auto-guidance, which employs a simplistic interpolation approach using a degraded model version to refine latent variables. While they shares a similar high-level concept, W2SD introduces fundamentally distinct mechanisms, delivering significantly stronger performance and broader applicability. Comprehensive quantitative and qualitative analyses are provided in Section E.4.
@@ 1-D Gaussian Mixture Data
Figure 4: Denoising trajectories across different settings (1-D Gauss). The weak model (blue) generates only right-peak data due to missing left-peak training samples, while the strong model (red) produces data between both peaks. W2SD balances the distribution by leveraging the reflective operator $\mathcal{M}_{\mathrm{inv}}^{\mathrm{w}}(\mathcal{M}^{\mathrm{s}}(\cdot))$. Additional examples are provided in  Figure 18.
@@ 2-D Gaussian Mixture Data
We also visualize the mechanism of W2SD on 2D scenario. In Figure 5 (first row), we modulate the proportion of the training data to obtain $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$. W2SD balances the distribution by exploiting the discrepancy between $\mathcal{M}^{\mathrm{s}}$ and $\mathcal{M}^{\mathrm{w}}$ (the region in the bottom-left corner, denoted as $\Delta_{1}$) to approximate the unattainable strong-to-ideal difference $\Delta_{2}$, and boost the chances of sampling toward the bottom-left region. In Figure 5 (second row), we visualize the denoising trajectories under different settings, further validating the effectiveness of W2SD.
@@ 6 Conclusion
To the best of our knowledge, this work is the first to systematically integrate the weak-to-strong concept into the inference enhancement of diffusion models via a reflection mechanism. We theoretically and empirically understand that the estimated weak-to-strong difference can effectively bridge the strong-to-ideal difference, enhancing the alignment between the learned distributions from existing diffusion models and the real data distribution. Building on this concept, we propose W2SD, which utilizes the estimated difference in density gradients to optimize sampling trajectories via reflective operations. W2SD demonstrates its effectiveness as a general-purpose framework through its cumulative performance improvements, flexible definition of weak-to-strong model pairs, and efficient performance gains with minimal computational overhead. We hope this work inspires interest in diffusion model scaling laws, advances sampling methods, and deepens understanding of probabilistic modeling in generative models.
@@ Appendix A Related Work
The concept of improving weak models into strong models originates from the AdaBoost (Høgsgaard et al., 2023), which constructs a more accurate classifier by aggregating multiple weak classifiers. Building upon this,  Green Larsen & Ritzert (2022); Høgsgaard et al. (2023) introduced a provably optimal weak-to-strong learner, establishing a robust theoretical foundation for this weak-to strong paradigm. In the field of LLMs, several studies (Chen et al., ; Burns et al., 2023) have utilized weak models as supervisory signals to facilitate the alignment of LLMs. This paradigm of weak-to-strong generation during training has similarly been investigated in the context of diffusion model training (Chen et al., 2025). And Sugiyama et al. (2013) recommended directly estimating the density difference between weak and strong models instead of separate estimations. And Karras et al. (2024) examines the mechanism of CFG and achieves improved results through interpolation-based perturbation. In contrast, W2SD utilizes the reflection by leveraging the gap between denoising and inversion to steer generation towards user-defined directions. And we note that W2SD generalizes the concept of ideal model, with experiments across human preference, personalization, fidelity and so on, confirming its wide applicability.
@@ D.1 Proof of Theorem 1
where $s_{\theta}^{\mathrm{s}}$ denotes the score predicted by $\mathcal{M}^{\mathrm{s}}$, which is equivalent to the gradient of the log probability density $\log{p_{t}^{\mathrm{s}}}$, so, Equation 7 can be rewritten as
$x_{t-\Delta t}=x_{t}+\sigma^{2t}\nabla_{x_{t}}\log{p_{t}^{\mathrm{s}}(x_{t})}\Delta t.$ (8)
@@
where $s_{\theta}^{\mathrm{w}}$ denotes the score predicted by $\mathcal{M}^{\mathrm{w}}$. Since $\Delta t$ is typically small, we neglect the approximation error in the diffusion inversion process which implies that $s_{\theta}^{\mathrm{w}}(x_{t-\Delta t},t)$ is equivalent to the gradient of the log probability density $\log{p_{t}^{\mathrm{w}}}$. Hence we can reformulate Equation 9 as
$\tilde{x}_{t}=x_{t-\Delta t}-\sigma^{2t}\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}\Delta t.$ (10)
@@
$\tilde{x}_{t}$ $=x_{t-\Delta t}-\sigma^{2t}\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}\Delta t$ (11)
$=x_{t}+\sigma^{2t}\nabla_{x_{t}}\log{p_{t}^{s}(x_{t})}\Delta t-\sigma^{2t}\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}\Delta t$ (12)
$=x_{t}+\sigma^{2t}\Delta t(\underbrace{\nabla_{x_{t}}\log{p_{t}^{s}(x_{t})}-\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}}_{\text{weak-to-strong difference}\Delta_{1}(t)}).$ (13)
From Equation 13, we observe that for the latent variable $x_{t}$, the reflection operator $\mathcal{M}^{\mathrm{w}}_{\mathrm{inv}}(\mathcal{M}^{\mathrm{s}}(\cdot))$ in W2SD perturbs $x_{t}$ along the direction of $\Delta_{1}(t)$, producing the refined variable $\tilde{x}_{t}$. When the weak-to-strong difference $\Delta_{1}$ closely bridges the unattainable strong-to-ideal difference $\Delta_{2}(t)$, the resulting $\tilde{x}_{t}$ becomes more aligned with the ground truth ideal distribution. ∎
@@ Appendix E Further analytical exploration
E.1 Deeper Visual Understanding
To provide a more intuitive demonstration of the W2SD effect, we present more 1D visualization examples in Figure 18 to aid understanding. As can be seen, W2SD has a certain probability of “correcting” the denoising trajectory of the model, resulting in more balanced generation outcomes. The settings here are consistent with those in Figure 4.
We also empirically validate that the weak-to-strong difference $\Delta_{2}$ can effectively bridges the strong-to-ideal difference $\Delta_{1}$. We first compute $\mathbf{CosineSimilarity}(\Delta_{1}(t),\Delta_{2}(t))$ at each timestep $t$. We train three models on CIFAR-10: the weak model (100 epochs), the strong model (300 epochs), and the ideal model (fully converged, 500 epochs). Here $\Delta_{1}(t)=\epsilon_{ideal}(t)-\epsilon_{strong}(t)$ and $\Delta_{1}(t)=\epsilon_{strong}(t)-\epsilon_{weak}(t)$. In the table below, the angle between $\Delta_{1}(t)$ and $\Delta_{2}(t)$ remains below $90^{o}$, confirming that weak-to-strong gap can reliably bridge strong-to-ideal gap.
Figure 19: Positive cosine similarity between $\Delta_{1}$ and $\Delta_{2}$ ($\cos{\Theta}<90$°).
To further validate this claim that $\Delta_{2}$ can bridge $\Delta_{1}$, we provide additional visualizations. As illustrated in Figure 20, the W2SD trajectory is visualized in 2D space, where its denoising direction (black) demonstrates closer alignment with the ideal direction (green). Moreover, the angles between the strong (red)-to-weak (blue) and ideal (green)-to-strong (red) vectors consistently remain within 90 °.
Figure 20: 2D denoising path visualization. The weak-to-strong gap can effectively bridge the strong-to-ideal gap.
E.2 The Impact of Approximation Error in the Inversion Process
@@
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}\log{p_{t}^{\mathrm{s}}(x_{t})}-\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}-\textrm{E}_{t}),$ (16)
@@
$\tilde{x}_{t}=x_{t}+\sigma^{2t}\Delta t(\nabla_{x_{t}}\log{p_{t}^{\mathrm{s}}(x_{t})}-\nabla_{x_{t}}\log{p_{t}^{\mathrm{w}}(x_{t})}-k\epsilon).$ (17)
@@
E.3 Relationship with Re-Sampling